OpenStudy (anonymous):

The perimeter of a rectangle is 20 cm. The width of the rectangle is 4 cm. What is the area of the rectangle? 4 cm2 8 cm2 12 cm2 24 cm2

OpenStudy (disco619):



OpenStudy (anonymous):

8

OpenStudy (shamil98):

No, the length of the rectangle is not equal to its width.

OpenStudy (disco619):

So we have two widths of the rectangle, which means we will remove both the widths of the rectangle to find the two lengths... = 20 - 4+4 = 20 - 8 = 12 Both lengths equal to 12cm, but we need only 1 length so we divide by two. \[1 length = \frac{ 12 }{ 2 }\] \[= 6\] 1 length of Rectangle = 6cm 1 width of Rectangle = 4cm Area = 6x4 Area = 24 square cm [Answer]

OpenStudy (shamil98):

4 = width P = 2L +2W 20 = 2L + 8 12 = 2L L = 6

OpenStudy (shamil98):

A = L * W , A = 6 * 4 a = 24

OpenStudy (disco619):

@shamil98 it's bad manners to interrupt someone else's work.

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

THANKS @Disco619

OpenStudy (anonymous):

AND @shamil98